### v1

# The Generation Gap: Exploring Age Bias in Large Language Models

Abstract
In this paper, we explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories. Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics. Additionally, we explore the impact of incorporating age identity information in prompts and observe challenges in mitigating value discrepancies with different age cohorts. Our findings highlight the age bias in LLMs and provide insights for future work.
The Generation Gap: Exploring Age Bias in Large Language Models
Siyang Liu1   Trish Maturi1   Siqi Shen1   Rada Mihalcea1 1 The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
1 Introduction
Widely used Large Language Models (LLMs) should be reflective of all age groups Dwivedi et al. (2021); Wang et al. (2019); Hong et al. (2023). Age statistics estimate that by 2030, 44.8% of the US population will be over 45 years old Vespa et al. (2018), and one in six people worldwide will be aged 60 years or over World Health Organization (2022). Analyzing how the values (e.g, religious values) in LLMs align with different age groups can enhance our understanding of the experience that users of different ages have with an LLM. For instance, for an older group that may exhibit less inclination towards new technologies Czaja et al. (2006); Colley and Comber (2003), an LLM that embodies the values of a tech-savvy individual may lead to less empathetic interactions. Minimizing the value disparities between LLMs and the older population has the potential to lead to better communication between these demographics and the digital products they engage with.
Figure 1: Age-related bias in LLMs on thirteen human value categories. Human values in this figure refer in particular to the US groups. Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM values and human values as age increases. A positive trend coefficient signifies the widening gap observed from younger to older age groups, thus indicating a model leaning towards younger age groups.
In this paper, we investigate whether and which values in LLMs are more aligned with specific age groups. Specifically, by using the World Value Survey Haerpfer et al. (2020), we prompt various LLMs to elicit their values on thirteen categories, employing eight format variations in prompts for robust testing. We observe a general inclination of LLM values towards younger demographics, as shown in Fig 1. We also demonstrate the specific categories of value and example inquiries where LLMs exhibit such age preferences (See Sec 4).
Furthermore, we study the effect of adding age identity information when prompting LLMs. Specifically, we instruct LLMs to use an age and country identity before requesting their responses. Surprisingly, we find that adding age identity fails to eliminate the value discrepancies with targeted age groups on eight out of thirteen categories (see Fig 4), despite occasional success in specific instances (See Sec 5).
We advocate for a heightened awareness within the research community regarding the potential age bias inherent in LLMs, particularly concerning their predisposition towards certain values. We also emphasize the complexities involved in calibrating prompt engineering to effectively address this bias.
2 Social Biases in LLMs
Recent advancements in LLMs have led to human-level or even surpassing performance across various NLP tasks Brown et al. (2020); Radford et al. (2019); Ouyang et al. (2022). However, there is a growing concern regarding the presence of social bias in these models Kasneci et al. (2023), especially as certain biases could result in discrimination and emotional harm to the impacted users. Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals, while struggling with predictive capabilities concerning Black and Asian groups Sun et al. (2023). Additionally, recent studies by Santurkar et al.; McGee; Atari et al. consistently highlight a left-leaning or democratic political inclination among LLM models. Despite extensive scrutiny on LLM bias Santurkar et al. (2023); Sun et al. (2023); Duan et al. (2024), the age-related preferences of LLMs remain less explored, and call for a comprehensive investigation for a more equitable and inclusive technological landscape.
Figure 2: Alignment rank of values of ChatGPT over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with ChatGPT in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
3 Analytic Method
3.1 Dataset Processing
We derive human values across age demographics and create prompts utilizing data from the 7th wave of the World Values Survey (WVS)  [^1] Statisitcs of WVS is accessible via Document-Online analysis. Haerpfer et al. (2020). The survey systematically probes 94278 individuals globally on thirteen categories, covering a range of social, political, economic, religious, and cultural values. To assess human values, we group the respondents by age group  [^2] Age groups are recorded as 18-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country. Subsequently, we compute the average values for each age group and country to represent their respective cohorts.
3.2 Prompting
Models.
We conduct our analysis on four LLMs: ChatGPT (GPT-3.5-turbo 0613), InstructGPT (GPT-3.5-turbo-instruct) OpenAI (2023), FLAN-T5XXL Chung et al. (2022), and FLAN-UL2 Tay (2023), spanning both open-source and close-source, chat-based and completion-based, as well as decoder-only and encoder-decoder architecture LLMs.
Prompts.
We identify three key components for each inquiry in the survey: context, question ID&content, and options. In addition, we add format variation—also known as spurious variation—into prompts, as previous research Shu et al. (2023) suggests inconsistent performance in LLMs after receiving a minor prompt variation.
By altering both the format and the order of inquiry components, we build a set of eight distinct prompt per inquiry as shown in Tab 3. We utilize the average outcomes across multiple prompt variations for our analysis to make our probing method more robust. Due to the variety in LLM’s capability in following instructions, we encountered seven types of unexpected reply and present our coping methods for each, as summarized in Tab 1. We observe a high agreement among responses induced by different prompts. Specifically, though we observed LLM’s instability to prompt variations (See analysis in Appx B), for 95.5% of inquiries, more than half of the responses are centered on the same choice or its adjacent options.
Unexpected Reply Type Example Coping Method
returning null value { ”Q1”: null} map null into missing code -2
unprompted responses answer Q1 to Qn when only asking Qn-m to Qn keep the answers of asked questions
redundant texts ”Answer = {‘Q1’, 1}” extract the json result
substandard json Q1:‘1’ manually correct
incompelete answer on binary question In true/false inquiry, only mention {‘Q1’: 1} instead of {‘Q1’:1, ‘Q2’:0} manually complete
inconsistent redundancy {‘Q1’:1} {‘Q1’:2} pick the firstly-shown item
constraint violation being required to mention up to 5 from 10 items, however return a json with more than 5 positive numbers remove json format requirement, and ask for a reply in natural language; manually understand
refusing to reply As an artificial intelligence, I don’t have personal views or sentiments fill out with a missing code -2
Table 1: Unexpected reply summary and corresponding coping intervention
3.3 Measures
We represent values belonging to a certain category as a vector. Each question in the WVS questionnaire is treated as a dimension (See the number of questions for each category in Tab 2). Calculating the distance between two vectors upon the original dimensionality often leads to a sensitivity to the dimensions of outliers or trivial information. Instead, we utilize a principle component analysis (PCA) Tipping and Bishop (1999) on a range of value vectors for learning smooth representations. Specifically, we collect 372 different value vectors representing people across 62 countries and six age groups. After performing min-max normalization and normal standardization, we conduct a PCA to reduce the vectors to a dimensionality of two. Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value vector represented by the $i$th age group be $x_{i},y_{i}]$. We derive three metrics below for our further analysis:
Euclidean Distance, the distance between two value vectors.
$d_{i}=\sqrt{{(x_{LLM}-x_{i})}^{2}+{(y_{LLM}-y_{i})}^{2}},$
where $(x_{LLM},y_{LLM})$ represents values of LLM. Alignment Rank, the reverse rank of distance between LLM values and people across six age groups.
$r_{i}=argsort(argsort([d_{1},...,d_{6}]))[i]$
Trend Coeficient, the slope of the gap between LLM values with human across six age groups.
$r_{i}=\beta+\alpha i$
$\alpha=\arg\max_{\alpha}(\sum_{i=1}^{6}(r_{i}-(\beta+\alpha i))^{2})$
4 Aligning with Which Age on Which Values?
Trend Observation.
As shown in Fig 1, we observe a general inclination of four popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient. Fig 2 exemplifies this bias for ChatGPT, where the model tends to have a higher alignment rank on younger groups on the most categories, despite few exceptions (see the left case in Fig 3). The results on other LLMs and other countries can be found in Appx D and E.
Case Study.
In Fig 3, we show two representative prompts and their responses from ChatGPT and human groups, to illustrate sample values where ChatGPT exhibits a clear bias toward a specific age group.
Figure 3: Two WVS prompts and their responses from LLMs and humans (in purple).
Figure 4: Change of Euclidean distance after adding identity information. The compared data is from values of ChatGPT and humans from different age groups in the US.
5 The Effect of Adding Identity in Prompts
Prompt Adjustment.
To analyze if adding age identity in the prompt helps to align values of LLM with the targeted age groups, we adjust our prompts by adding a sentence like “Suppose you are from [country] and your age is between [lowerbound] and [upperbound].” at the beginning of the required component of the original prompt and get responses that corresponds with six age groups.
Observation on Gap Change.
We illustrate the change of Euclidean distance between values of LLM and different age groups after adding identity information. As is presented in Fig 4, in eight out of thirteen categories (No.1,2,4,5,7,8,9,12) no improvement is observed.
Case Study.
We also showcase a successful calibration example for a question from the political interest WVS section in Fig 5. The value pyramid illustrates LLMs’ responses for different age ranges compares to the answers from the U.S. population. When age is factored into the LLM prompt, the LLM’s views are more aligned with the U.S. population of that respective age group, as it reports higher frequency using radio news for the older group.
6 Recommendations for Future Work
We have observed that simply including an age in prompts fails to eliminate the value disparity for the targeted age groups. Out of the thirteen categories inquired upon, eight have shown no improvement. To this end, we recommend a careful data curation during pretraining. Doing so involves a deliberate and thoughtful selection of data sources that are diverse and representative of various age groups. By doing so, we can ensure that the model’s training material reflects a wide range of perspectives and experiences, thereby reducing biases and disparities in the model’s responses. We also recommend a consideration of human feedback optimization (e.g., RLHF). Through this iterative process, LLMs can learn to generate responses that fit better with the needs of different age groups. These strategies help mitigate the value disparities associated with targeted age groups, enhancing the LLM’s abilities to be more equitable and inclusive.
Figure 5: Value Pyramid of U.S population (left) and ChatGPT (right) for an inquiry on the frequency of using radio news.
7 Conclusion
In this paper, we investigated the alignment of values in LLMs with specific age groups using data from the World Value Survey. Our findings suggest a general inclination of LLM values towards younger demographics. Our study contributes to raising attention to the potential age bias in LLMs and advocate continued efforts from the community to address this issue. Moving forward, efforts to calibrate value inclinations in LLMs should consider the complexities involved in prompting engineering and strive for equitable representation across diverse age cohorts.
Limitations
There are several limitations in our paper. Firstly, due to the time and cost, we were not able to try more sophisticated prompts for the age alignment, which may effectively eliminate the value disparity with targeted age groups. Secondly, our analysis relies on the questionnaire of WVS. However, their question design is not perfectly tailored for characterizing age discrepancies, which limits the depth of sights we could get from analysis. Finally, the range of LLMs in our analysis could be expanded.
Ethics Statement
Several ethical considerations have been included thorough our projects. Firstly, the acquisition of WVS data is under the permission of data publisher. Secondly, we carefully present our data analysis results with an academic honesty. This project is under a collaboration, we well-acknowledge the work of each contributor and ensure a transparent and ethical process thorough the whole collaboration. Finally, we leverage the ability of AI-assistants to help with improving paper writing while we guarantee the originality of paper content and have reviewed the paper by every word.
Appendix A World Value Survey
The WVS [^3] https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values. See the statistics of inquiries in Fig 2. Note that we remove ten of them that requires demographic information, as these are impossible for applying to an LLM lacking demographic data, and keep 249 inquiries as our final choices for prompting.
Value Category # Inquiry Example
Social Values, Norm, Stereotypes 45 how important family is in your life? (1:Very important, 2:Rather important, 3:Not very important, 4: Not at all important)
Happiness and Wellbeing 11 taking all things together, would you say you are? (1:1:Very happy, 2:Rather happy, 3:Not very happy, 4:Not at all happy)
Social Capital, Trust and Organizational Membership 49 would you say that most people can be trusted or that you need to be very careful in dealing with people? (1:Most people can be trusted, 2:Need to be very careful)
Economic Values 6 Which of them comes closer to your own point of view? (1:Protecting the environment should be given priority, even if it causes slower economic growth and some loss of jobs, 2:Economic growth and creating jobs should be the top priority, even if the environment suffers to some extent, 3:Other answer)
Perceptions of Migration 10 how would you evaluate the impact of these people on the development of your country? (1:Very good, 2:Quite good, 3:Neither good, nor bad, 4:Quite bad, 5:Very bad)
Perceptions of Security 21 could you tell me how secure do you feel these days? (1: Very secure, 2: Quite secure, 3: Not very secure, 4: Not at all secure)
Perceptions of Corruption 9 tell me for people in state authorities if you believe it is none of them, few of them, most of them or all of them are involved in corruption? (1:None of them, 2:Few of them, 3:Most of them, 4:All of them)
Index of Postmaterialism 6 if you had to choose, which of the following statements would you say is the most important? (1: Maintaining order in the nation, 2: Giving people more say in important government decisions, 3: Fighting rising prices, 4: Protecting freedom of speech,)
Perceptions about Science and Technology 6 it is not important for me to know about science in my daily life. (1:Completely disagree, 2:Completely agree)
Religious Values 8 The only acceptable religion is my religion (1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree)
Ethical Values 13 Abortion is? (1: Never justifiable, 10: Always justifiable)
Political Interest and Political Participation 36 Election officials are fair. (1:Very often,2:Fairly often,3:Not often,4:Not at all often)
Political Culture and Political Regimes 25 How important is it for you to live in a country that is governed democratically? On this scale where 1 means it is “not at all important” and 10 means “absolutely important” what position would you choose? (1:Not at all important, 10:Absolutely important)
Table 2: Statistics of inquires in World Value Survey.
Component Variant ID Example
Context 1 I’d like to ask you how much you trust people from various groups. Could you tell me for each whether you trust people from this group completely, somewhat, not very much or not at all?
QID and Content Unique ID 2.1 Q58: Your family Q59: Your neighborhood
Relative ID 2.2 Q1: Your family Q2: Your neighborhood
Options Style1 3.1 Options: 1:Trust completely, 2:Trust somewhat, 3:Do not trust very much, 4:Do not trust at all
Style2 3.2 Options: 1 represents Trust completely, 2 represents Trust somewhat, 3 represents Do not trust very much, 4 represents Do not trust at all
Requirement Chat 4.1 Answer in JSON format, where the key should be a string of the question id (e.g., Q1), and the value should be an integer of the answer id.
Completion 4.2 Answer in JSON format, where the key should be a string of the question id (e.g., Q1), and the value should be an integer of the answer id. The answer is
(a) Inquiry Components and Corresponding Prompt Variants
Order of Prompt
1  2.13.14.x
1 2.23.14.x
13.12.14.x
13.12.24.x
12.13.24.x
12.23.24.x
13.22.14.x
13.22.24.x
(b) Prompt Orders
Table 3: Prompt Pipeline Details
Appendix B The Instability of LLM Outputs Due to Prompt Variations
Regarding the unstableness of LLM outputs due to prompting variation, we observed LLM’s instability to prompt variations. However, instead of testing more prompts, we ended up using the designed eight variations to support our study. Our decision was made by conducting a deep analysis of using our current prompts. The key findings are listed below:
  1. (1)
    56.3% of survey questions exhibited inconsistent answers induced by eight different prompts.
  2. (2)
    In 68.1% of survey questions , six or more prompts resulted in the majority answer.
  3. (3)
    In 80.3% of survey questions, four or more prompts induce the majority answer.
  4. (4)
    For 45 questions, fewer than four prompts led to the majority answer, indicating diverse choices and reflecting LLMs’ self-conflict on these questions. These questions are on economic equity/liberty, sex conservation/freedom, whether acknowledging the importance of developing economics, perception about the living environment, etc.
  5. (5)
    Despite potential variations in answers induced by prompt variation, we found for 95.5% of inquiries, more than half of the responses are centered on the same choice or its adjacent options. Adjacent option is a score equal to the majority score +/- 1.
Eventually, while discovering the unstableness of LLM outputs, we believe it is reasonable to use the average score from eight prompts as a representative value.
Appendix C Prompting Details
Our prompting process can be described as three steps below:
  1. 1.
    Repeatedly request LLMs’ responses on survey questions with 8 different prompts. For each question, there will be 8 numerical scores induced by prompts,where only the missing code is a negative number.
  2. 2.
    Calculate the mean of scores for each question while ignoring negative scores. Then we can get vectors that consist of scores from questions for each value category. The vector represents the LLM’s value in a specific category.
  3. 3.
    Preprocess the value vector for data analysis, as illustrated in Sec 3.1.
The cost of API calling from Closed-coursed LLMs is less than 5 dollars. For the deployment of FLAN-T5-XXL and FLAN-UL2 models, we ran either model on a single A40 GPU with float16 precision. When prompting, we prompt models with a temperature 1.0, max token length 1024, random seed 43.
Appendix D Results on Other LLMs
In the section, we supplement the alignment ranking results on InstructGPT (Fig 6), FLAN-T5-XXL (Fig 7) and FLAN-UL2 (Fig 8) respectively.
Appendix E Results on Other Countries
We have extended our analysis to include alignment results from an additional four pairs of countries: Argentina and Brazil (Fig 4), Ethiopia and Nigeria (Fig 5), Germany and Great Britain (Fig 6), and Indonesia and Malaysia (Fig 7) .
Figure 6: Alignment rank of values of InstructGPT over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with InstructGPT in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 7: Alignment rank of values of FLAN-T5-XXL over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with FLAN-T5-XXL in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Figure 8: Alignment rank of values of FLAN-UL2 over different age groups in the US. Rank 1 on an specific age group represents that this age group has the narrowest gap with FLAN-UL2 in values. A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
Table 4: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
Table 5: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
Table 6: Alignment rank of LLMs over different age groups in Gemany and Great Britain. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
Table 7: Alignment rank of LLMs over different age groups in Indonesia and Malaysia. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT

### v2

@@
-The Generation Gap: Exploring Age Bias in Large Language Models
+The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models
 Abstract
 In this paper, we explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories.
-Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics.
+Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics, especially in the US.
 Additionally, we explore the impact of incorporating age identity information in prompts and observe challenges in mitigating value discrepancies with different age cohorts.
 Our findings highlight the age bias in LLMs and provide insights for future work.
-The Generation Gap: Exploring Age Bias in Large Language Models
-Siyang Liu1   Trish Maturi1   Siqi Shen1   Rada Mihalcea1 1 The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
+Materials for our analysis will be available via anonymous.github.com
+The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models
+Siyang Liu1   Trish Maturi1   Bowen Yi1   Siqi Shen1   Rada Mihalcea1 1 The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
 1 Introduction
@@
 Human values in this figure refer in particular to the US groups.
-Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM values and human values as age increases.
-A positive trend coefficient signifies the widening gap observed from younger to older age groups, thus indicating a model leaning towards younger age groups.
+Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM and human values as age increases.
+A positive trend coefficient signifies the widening gap observed from younger to older groups, thus indicating a model leaning towards younger age groups.
+Significant test is detailed in Appx G
 In this paper, we investigate whether and which values in LLMs are more aligned with specific age groups.
@@
 Surprisingly, we find that adding age identity fails to eliminate the value discrepancies with targeted age groups on eight out of thirteen categories (see Fig 4), despite occasional success in specific instances (See Sec 5).
-We advocate for a heightened awareness within the research community regarding the potential age bias inherent in LLMs, particularly concerning their predisposition towards certain values.
-We also emphasize the complexities involved in calibrating prompt engineering to effectively address this bias.
-2 Social Biases in LLMs
-Recent advancements in LLMs have led to human-level or even surpassing performance across various NLP tasks Brown et al.
+We advocate for increased awareness within the research community regarding the potential age bias inherent in LLMs, particularly concerning their predisposition towards certain values.
+We also emphasize the complexities involved in calibrating prompts to effectively address this bias.
+2 Related Work
+Due to the recent advancements in LLMs in manifesting human-level performance across various tasks Brown et al.
 (2020); Radford et al.
 (2019); Ouyang et al.
+(2022), there is a growing concern regarding the presence of social bias in these models Kasneci et al.
 (2023).
-However, there is a growing concern regarding the presence of social bias in these models Kasneci et al.
-(2023), especially as certain biases could result in discrimination and emotional harm to the impacted users.
-Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals, while struggling with predictive capabilities concerning Black and Asian groups Sun et al.
+Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals Sun et al.
+(2023), and political inclination Santurkar et al.
+(2023); McGee (2023); Atari et al.
 (2023).
-Additionally, recent studies by Santurkar et al.; McGee; Atari et al.
-consistently highlight a left-leaning or democratic political inclination among LLM models.
 Despite extensive scrutiny on LLM bias Santurkar et al.
 (2023); Sun et al.
-(2023); Duan et al.
-(2024), the age-related preferences of LLMs remain less explored, and call for a comprehensive investigation for a more equitable and inclusive technological landscape.
-Figure 2: Alignment rank of values of ChatGPT over different age groups in the US.
-Rank 1 on an specific age group represents that this age group has the narrowest gap with ChatGPT in values.
-A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
+(2023), the age-related preferences of LLMs remain less explored.
+Previous work has mentioned age as one of multi-facets of bias in LLM performance Kamruzzaman et al.
+(2023); Haller et al.
+(2023); Draxler et al.
+(2023); Levy et al.
+(2024); Oketunji et al.
+(2023), while lacking a direct study on the age aspect.
+Recent research Duan et al.
+(2024) publishes an evaluation for well-known LLMs on age bias through 50 multi-choice questions; unlike it focuses on discriminatory narratives towards specific age groups, our investigation is running at an implicit level.
+We argue that understanding the underlying value systems is crucial, as the value discrepancies between users and LLMs can significantly impact their adoption of LLMs, even though the explicit discrimination is rectified, as exemplified in our discussion on technology attitudes in Sec 1.
+Figure 2: Alignment rank of values of LLMs over different age groups in specific Countries.
+See results on more models and countries in Appx E and F .
+Rank 1 on an specific age group represents that this age group has the narrowest gap with LLM in values.
+A increasing monoticity indicates a closer alignment towards younger groups.
 3 Analytic Method
-3.1 Dataset Processing
-We derive human values across age demographics and create prompts utilizing data from the 7th wave of the World Values Survey (WVS)  [^1] Statisitcs of WVS is accessible via Document-Online analysis.
-Haerpfer et al.
+3.1 Human Data Acquisition
+Dataset.
+We derive human values utilizing the 7th wave of the World Values Survey (WVS) Haerpfer et al.
 (2020).
-The survey systematically probes 94278 individuals globally on thirteen categories, covering a range of social, political, economic, religious, and cultural values.
+The survey systematically probes 94k individuals globally on 13 categories, covering a range of social, political, economic, religious, and cultural values.
+See an introduction of WVS in Appx A.
+Each inquiry is a single-choice question.
+Responses are numeric, quantifying the inclination on the options, e.g., “1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree”.
+Negative number is possible for coding exceptions such as “I don’t know”.
 To assess human values, we group the respondents by age group  [^1] Age groups are recorded as 18-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country.
-Subsequently, we compute the average values for each age group and country to represent their respective cohorts.
+Subsequently, we compute the average values for each age group and country to represent their respective cohorts, ignoring the invalid negative numbers.
 3.2 Prompting
 Models.
-We conduct our analysis on four LLMs: ChatGPT (GPT-3.5-turbo 0613), InstructGPT (GPT-3.5-turbo-instruct) OpenAI (2023), FLAN-T5XXL Chung et al.
-(2022), and FLAN-UL2 Tay (2023), spanning both open-source and close-source, chat-based and completion-based, as well as decoder-only and encoder-decoder architecture LLMs.
+We conduct our analysis on six LLMs, as introduced in Tab 1.
+Model (Version) Features
+ChatGPT(GPT-3.5-turbo 0613)
+InstructGPT (GPT-3.5-turbo-instruct)
+Mistral (mistral-7B-v0.1)
+Vicuna (vicuna-7b-v1.5)
+FLAN-T5 (flan-t5-xxl)
+FLAN-UL2 (flan-ul2)
+Table 1: Model description.
+: commercial models,
+: open models,
+: chat-based,
+: completion-based,
+: RLHF, and
+: training with instructions.
 Prompts.
 We identify three key components for each inquiry in the survey: context, question ID&content, and options.
-In addition, we add format variation—also known as spurious variation—into prompts, as previous research Shu et al.
-(2023) suggests inconsistent performance in LLMs after receiving a minor prompt variation.
-By altering both the format and the order of inquiry components, we build a set of eight distinct prompt per inquiry as shown in Tab 3.
-We utilize the average outcomes across multiple prompt variations for our analysis to make our probing method more robust.
-Due to the variety in LLM’s capability in following instructions, we encountered seven types of unexpected reply and present our coping methods for each, as summarized in Tab 1.
-We observe a high agreement among responses induced by different prompts.
-Specifically, though we observed LLM’s instability to prompt variations (See analysis in Appx B), for 95.5% of inquiries, more than half of the responses are centered on the same choice or its adjacent options.
-Unexpected Reply Type Example Coping Method
-returning null value { ”Q1”: null} map null into missing code -2
-unprompted responses answer Q1 to Qn when only asking Qn-m to Qn keep the answers of asked questions
-redundant texts ”Answer = {‘Q1’, 1}” extract the json result
-substandard json Q1:‘1’ manually correct
-incompelete answer on binary question In true/false inquiry, only mention {‘Q1’: 1} instead of {‘Q1’:1, ‘Q2’:0} manually complete
-inconsistent redundancy {‘Q1’:1} {‘Q1’:2} pick the firstly-shown item
-constraint violation being required to mention up to 5 from 10 items, however return a json with more than 5 positive numbers remove json format requirement, and ask for a reply in natural language; manually understand
-refusing to reply As an artificial intelligence, I don’t have personal views or sentiments fill out with a missing code -2
-Table 1: Unexpected reply summary and corresponding coping intervention
+To ensure robustness, we made several format variations for the prompt [^2] Despite adopting format variations, we were cautious to not include major changes as the content and structure of WVS were carefully designed by sociologists and professionals.
+(e.g., alter wordings and change order of components), as previous research Shu et al.
+(2023); Röttger et al.
+(2024); Beck et al.
+(2023) uncovered inconsistent performance in LLMs after receiving a minor prompt variation.
+Eventually, we build a set of eight distinct prompts per inquiry.
+Please see prompt design details in Tab 3 in Appendix.
+Through a careful analysis on the prompt responses (Appx B), We observe unstableness of LLM’s responses to prompt variations.
+However, multiple prompt trials assists with achieving a convergence point.
+On 95.5% of questions, more than half of the eight prompts led to responses centered on the same choice or adjacent options, and thus we believe it is acceptable to consider the average of the outcomes across the eight prompt variations as the LLM’s final responses to WVS.
+In addition, due to the instability of LLMs in following instructions, we encountered seven types of unexpected reply and present our coping methods for each, as summarized in Tab 4.
+In the process of averaging responses, we ignore the invalid negative numbers, as we did in calculating human values.
+For reproducing our work, parameter setting and prompting details are reported in Appendix D.
 3.3 Measures
-We represent values belonging to a certain category as a vector.
-Each question in the WVS questionnaire is treated as a dimension (See the number of questions for each category in Tab 2).
-Calculating the distance between two vectors upon the original dimensionality often leads to a sensitivity to the dimensions of outliers or trivial information.
-Instead, we utilize a principle component analysis (PCA) Tipping and Bishop (1999) on a range of value vectors for learning smooth representations.
-Specifically, we collect 372 different value vectors representing people across 62 countries and six age groups.
-After performing min-max normalization and normal standardization, we conduct a PCA to reduce the vectors to a dimensionality of two.
-Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value vector represented by the $i$th age group be $x_{i},y_{i}]$.
-We derive three metrics below for our further analysis:
-Euclidean Distance, the distance between two value vectors.
-$d_{i}=\sqrt{{(x_{LLM}-x_{i})}^{2}+{(y_{LLM}-y_{i})}^{2}},$
-where $(x_{LLM},y_{LLM})$ represents values of LLM. Alignment Rank, the reverse rank of distance between LLM values and people across six age groups.
-$r_{i}=argsort(argsort([d_{1},...,d_{6}]))[i]$
-Trend Coeficient, the slope of the gap between LLM values with human across six age groups.
-$r_{i}=\beta+\alpha i$
-$\alpha=\arg\max_{\alpha}(\sum_{i=1}^{6}(r_{i}-(\beta+\alpha i))^{2})$
+We use vector $V_{c}$ to represent values belonging to a certain category $c$. Each question in the WVS questionnaire is treated as a dimension:
+[…]
+where $r_{i}$ is a numeric response to the $i$th question in the section of $c$, and $n_{c}$ denotes the total question number.
+Note the acquisition of numeric responses for human groups and LLM has been illustrated in Sec 3.1 and 3.2.
+By collecting 372 value vectors that represent people across 62 countries and 6 age groups, along with a value vector for the LLM to compare, we utilize principle component analysis (PCA) Tipping and Bishop (1999) on totally 373 value vectors for representation learning.
+We acquire value representations for all groups with the dimensionality of three.
+Our consideration of using PCA is added in Appx C.
+[…]
+Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value representation for the $i$th age group be $x_{c,i},y_{c,i},z_{c,i}]$.
+We derive three metrics below for our further analyses:
+Euclidean Distance, the distance between two value representations.
+[…]
+where $(x_{c,M},y_{c,M},z_{c,M})$ represents values of LLM on category $c$.
+Alignment Rank, the ascending rank of distances between LLM values and people across six age groups.
+[…]
+Trend Coefficient, the slope of the value gap between LLM and humans across six age groups. $\alpha$ is the slope we would like to fit by linear regression.
+[…]
+[…]
 4 Aligning with Which Age on Which Values?
 Trend Observation.
-As shown in Fig 1, we observe a general inclination of four popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient.
-Fig 2 exemplifies this bias for ChatGPT, where the model tends to have a higher alignment rank on younger groups on the most categories, despite few exceptions (see the left case in Fig 3).
-The results on other LLMs and other countries can be found in Appx D and E.
+As shown in Fig 1, we observe a general inclination of popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient.
+Fig 2 exemplifies the bias for LMMs across six age groups in several countries.
+Due to the limited paper pages, results on other LLMs and countries can be found in Appx E and F.
+Significant testing procedure is available in Appx G.
+We observe that in the US and China, as countries of large population, the models tend to have a higher alignment rank on younger groups on the most categories, despite few exceptions (e.g., happiness and well-being).
+However, in Ethiopia and Nigeria (Tab 8), the inclination is less evident.
+We leave this phenomenon for future study.
+Figure 3: Two WVS prompts and their responses from LLMs and humans (in purple).
 Case Study.
 In Fig 3, we show two representative prompts and their responses from ChatGPT and human groups, to illustrate sample values where ChatGPT exhibits a clear bias toward a specific age group.
-Figure 3: Two WVS prompts and their responses from LLMs and humans (in purple).
-Figure 4: Change of Euclidean distance after adding identity information. The compared data is from values of ChatGPT and humans from different age groups in the US.
 5 The Effect of Adding Identity in Prompts
@@
 As is presented in Fig 4, in eight out of thirteen categories (No.1,2,4,5,7,8,9,12) no improvement is observed.
+Figure 4: Change of Euclidean distance after adding identity information. The compared data is from values of ChatGPT and humans from different age groups in the US.
 Case Study.
-We also showcase a successful calibration example for a question from the political interest WVS section in Fig 5.
+We also showcase a successful calibration example for a question about the source of acquiring information in Fig 5.
 The value pyramid illustrates LLMs’ responses for different age ranges compares to the answers from the U.S.
@@
 population of that respective age group, as it reports higher frequency using radio news for the older group.
+Figure 5: Value Pyramid of U.S population (left) and ChatGPT (right) for an inquiry on the frequency of using radio news.
 6 Recommendations for Future Work
@@
 These strategies help mitigate the value disparities associated with targeted age groups, enhancing the LLM’s abilities to be more equitable and inclusive.
-Figure 5: Value Pyramid of U.S population (left) and ChatGPT (right) for an inquiry on the frequency of using radio news.
 7 Conclusion
@@
 The WVS [^3] https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values.
+We share a page of WVS questionnaire in Tab 6.
 See the statistics of inquiries in Fig 2.
+Demographic statistics of WVS is accessible via Document-Online analysis.
 Note that we remove ten of them that requires demographic information, as these are impossible for applying to an LLM lacking demographic data, and keep 249 inquiries as our final choices for prompting.
@@
 13.22.24.x
-(b) Prompt Orders
+(b) Eight Prompts with Changing Orders
+An Example Prompt for Order 1 2.2 3.1 4.1
+For each of the following statements I read out, can you tell me how strongly you agree or disagree with each.
+Do you strongly agree, agree, disagree, or strongly disagree?
+Q1:One of my main goals in life has been to make my parents proud.
+Options: 1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree.
+Answer in JSON format, where the key should be a string of the question id (e.g., Q1), and the value should be an integer of the answer id.
+(c) Example Prompt
 Table 3: Prompt Pipeline Details
+Unexpected Reply Type Example Coping Method
+returning null value { ”Q1”: null} map null into missing code -2
+unprompted responses answer Q1 to Qn when only asking Qn-m to Qn keep the answers of asked questions
+redundant texts ”Answer = {‘Q1’, 1}” extract the json result
+substandard json Q1:‘1’ manually correct
+incompelete answer on binary question In true/false inquiry, only mention {‘Q1’: 1} instead of {‘Q1’:1, ‘Q2’:0} manually complete
+inconsistent redundancy {‘Q1’:1} {‘Q1’:2} pick the firstly-shown item
+constraint violation being required to mention up to 5 from 10 items, however return a json with more than 5 positive numbers remove json format requirement, and ask for a reply in natural language; manually understand
+refusing to reply As an artificial intelligence, I don’t have personal views or sentiments fill out with a missing code -2
+Table 4: Unexpected reply summary and corresponding coping intervention
 Appendix B The Instability of LLM Outputs Due to Prompt Variations
@@
 Eventually, while discovering the unstableness of LLM outputs, we believe it is reasonable to use the average score from eight prompts as a representative value.
-Appendix C Prompting Details
+Appendix C Reasons of Applying PCA
+  1. 1.
+    Each question in WVS ought not to be equally important.
+Furthermore, for the questions belonging to a certain category, they correlate with each other.
+To this end, we need to find out the principle components among multi inquiries.
+  2. 2.
+    PCA here is also used as an unsupervised representation learning method.
+Compared to utilizing original data, the representations learned from hundreds of comparable examples (372 value vectors from different country and age groups) will mitigate the curse of dimensionality and other undesired properties of high-dimensional spaces.
+Other representation learning methods are also applicable.
+As the medium number of original dimensionality for all categories is 11, PCA is enough to handle the learning problem.
+Appendix D Prompting Details
 Our prompting process can be described as three steps below:
@@
 The cost of API calling from Closed-coursed LLMs is less than 5 dollars.
-For the deployment of FLAN-T5-XXL and FLAN-UL2 models, we ran either model on a single A40 GPU with float16 precision.
-When prompting, we prompt models with a temperature 1.0, max token length 1024, random seed 43.
-Appendix D Results on Other LLMs
+For the deployment of open-sourced models, we ran either model on a single A40 GPU with float16 precision.
+When prompting, we prompt models with a temperature 1.0, max token length 1024, random seed 42.
+Appendix E Results on Other LLMs
 In the section, we supplement the alignment ranking results on InstructGPT (Fig 7), FLAN-T5-XXL (Fig 8) and FLAN-UL2 (Fig 9) respectively.
-Appendix E Results on Other Countries
+Appendix F Results on Other Countries
 We have extended our analysis to include alignment results from an additional four pairs of countries: Argentina and Brazil (Fig 7), Ethiopia and Nigeria (Fig 8), Germany and Great Britain (Fig 9), and Indonesia and Malaysia (Fig 10).
+Appendix G Significant Test
+In this section, we conduct two kinds of significant tests to support our study: (1) we use MANOVA to test the significant difference among human values from different age groups, and (2) we use t-distribution to test the significant tendency of LLMs towards younger group.
+Notes our focus lies in characterizing the inclination of LLM values towards specific age groups.
+So, we are claiming significant tendency instead of claiming LLMs significantly assemble of any specific age groups.
+G.1 Significant Test for the Discrepancy among Human Age Groups
+Our analysis should base on a reasonable precondition that in WVS, human values significantly diverse cross different age groups.
+We used MANOVA (multivariate analysis of variance) to test the significant difference of human values cross all age groups, as shown below:
+Null hypothesis ($H_{0}$): the age group has no effect on any responses to the survey questions
+Statistics: Wilks’ lambda
+Result: See Tab 5. In conclusion: We reject the null hypothesis with p-value ¡ 1e-4
+Country Value Num DF Den DF F Value Pr ¿ F (p-value)
+US 0.07 176.00 1631.00 124.82 0.0000*
+China 0.06 184.00 2068.00 164.16 0.0000*
+Germany 0.05 118.00 1048.00 173.11 0.0000*
+Great British 0.06 118.00 1607.00 220.91 0.0000*
+Indonesia 0.09 201.00 2310.00 113.78 0.0000*
+Malaysia 0.09 254.00 1022.00 42.43 0.0000*
+Ethiopia 0.16 127.00 843.00 34.02 0.0000*
+Nigeria 0.13 176.00 614.00 23.18 0.0000*
+Table 5: P-values of value difference among different age groups in specific countries. * indicates p-value¡1e-4
+G.2 Significant Test for Trend Coefficient
+As it may be hard to interpret the trend coefficient in Fig 1 on some categories (e.g., perception of corruption).
+Despite its bias towards younger/older, it may not be a significantly meaningful number.
+We add significance testing for the linear regression on trend coefficient.
+Null hypothesis ($H_{0}$): $\alpha=0$, where is the trend coefficient fitted by a linear regression model presented in Sec 3.3.
+Statistics: t distribution.
+Results: see Tab 6.
+Category ChatGPT InstructGPT Mistral Vicuna Flan-t5 Flan-ul
+Social Values, Norm, Stereotypes 0.33 0.111 0.208 0.072* 0.005* 0.042*
+Happiness and Wellbeing 0.042* 0.208 0.005* 0.005* 0.005* 0.005*
+Social Capital, Trust and Organizational 0.397 0.872 0.005* 0.000* 0.042* 0.397
+Economic Values 0.000* 0.468 0.872 0.468 0.623 0.042*
+Perceptions of Corruption 0.704 0.072* 0.019* 0.072* 0.019* 0.005*
+Perceptions of Migration 0.072* 0.042* 0.005* 0.266 0.000* 0.156
+Perceptions of Security 0.042* 0.000* 0.000* 0.000* 0.000* 0.000*
+Index of Postmaterialism 0.623 0.787 0.397 0.111 0.787 0.005*
+Perceptions about Science and Technology 0.329 0.468 0.329 0.005* 0.329 0.623
+Religious Values 0.111 0.544 0.005* 0.005* 0.005* 0.019*
+Ethical Values 0.000* 0.000* 0.000* 0.000* 0.072* 0.000*
+Political Interest and Political Participation 0.208 0.872 0.000* 0.000* 0.208 0.329
+Political Culture and Political Regimes 0.000* 0.000* 0.000* 0.005* 0.957 0.872
+Table 6: P-values of trend coefficients for each model on each value category. * indicates p-value¡0.1
 Figure 7: Alignment rank of values of InstructGPT over different age groups in the US.
@@
 A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
-Table 4: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
-Table 5: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
-Table 6: Alignment rank of LLMs over different age groups in Gemany and Great Britain. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
-Table 7: Alignment rank of LLMs over different age groups in Indonesia and Malaysia. LLM tested in each image is (a) ChatGPT, (b)Flan-t5-xxl, (c)Flan-ul, and (d)InstructGPT
+Table 7: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
+Table 8: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
+Table 9: Alignment rank of LLMs over different age groups in Gemany and Great Britain.
+LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
+Table 10: Alignment rank of LLMs over different age groups in Indonesia and Malaysia.
+LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.

### v3

@@
-The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models
+The Generation Gap: Exploring Age Bias in the Value Systems of Large Language Models
 Abstract
-In this paper, we explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories.
-Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics, especially in the US.
+We explore the alignment of values in Large Language Models (LLMs) with specific age groups, leveraging data from the World Value Survey across thirteen categories.
+Through a diverse set of prompts tailored to ensure response robustness, we find a general inclination of LLM values towards younger demographics, especially when compared to the US population.
+Although a general inclination can be observed, we also found that this inclination toward younger groups can be different across different value categories.
 Additionally, we explore the impact of incorporating age identity information in prompts and observe challenges in mitigating value discrepancies with different age cohorts.
 Our findings highlight the age bias in LLMs and provide insights for future work.
-Materials for our analysis will be available via anonymous.github.com
-The Generation Gap: Exploring Age Bias Underlying in the Value Systems of Large Language Models
-Siyang Liu1   Trish Maturi1   Bowen Yi1   Siqi Shen1   Rada Mihalcea1 1 The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
+Materials for our analysis are available at https://github.com/MichiganNLP/Age-Bias-In-LLMs
+The Generation Gap: Exploring Age Bias in the Value Systems
+of Large Language Models
+Siyang Liu  Trisha Maturi  Bowen Yi  Siqi Shen  Rada Mihalcea The LIT Group, Department of Computer Science and Engineering, University of Michigan, Ann Arbor lsiyang@umich.edu, mihalcea@umich.edu
+Figure 1: Age-related bias in LLMs on thirteen human value categories.
+Human values in this figure refer in particular to the US groups.
+Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM and human values as age increases.
+A positive trend coefficient signifies the widening gap observed from younger to older groups, thus indicating a model leaning towards younger age groups.
+The significance test is detailed in Appx F.
 1 Introduction
@@
 (2018), and one in six people worldwide will be aged 60 years or over World Health Organization (2022).
-Analyzing how the values (e.g, religious values) in LLMs align with different age groups can enhance our understanding of the experience that users of different ages have with an LLM.
+Analyzing how the values (e.g., religious values) in LLMs align with different age groups can enhance our understanding of the experience that users of different ages have with an LLM.
 For instance, for an older group that may exhibit less inclination towards new technologies Czaja et al.
@@
 Minimizing the value disparities between LLMs and the older population has the potential to lead to better communication between these demographics and the digital products they engage with.
-Figure 1: Age-related bias in LLMs on thirteen human value categories.
-Human values in this figure refer in particular to the US groups.
-Trend coefficients (see calculation in Sec 3.3) were derived from the slope of the changing gap between LLM and human values as age increases.
-A positive trend coefficient signifies the widening gap observed from younger to older groups, thus indicating a model leaning towards younger age groups.
-Significant test is detailed in Appx G
 In this paper, we investigate whether and which values in LLMs are more aligned with specific age groups.
@@
 2 Related Work
-Due to the recent advancements in LLMs in manifesting human-level performance across various tasks Brown et al.
-(2020); Radford et al.
+Due to the rapid advancements in LLMs across various tasks Brown et al.
 (2020); Ouyang et al.
@@
 Recent research has shown that LLMs exhibit “preferences” for certain demographic groups, such as White and female individuals Sun et al.
-(2023), and political inclination Santurkar et al.
-(2023); McGee (2023); Atari et al.
+(2023), and political inclination McGee (2023); Atari et al.
 (2023).
-Despite extensive scrutiny on LLM bias Santurkar et al.
-(2023); Sun et al.
-(2023), the age-related preferences of LLMs remain less explored.
-Previous work has mentioned age as one of multi-facets of bias in LLM performance Kamruzzaman et al.
+However, the age-related preferences of LLMs remain less explored.
+Prior work has mentioned age as one of multi-facets of bias in LLM performance Kamruzzaman et al.
 (2023); Haller et al.
@@
 (2024); Oketunji et al.
-(2023), while lacking a direct study on the age aspect.
+(2023) while lacking a direct study on the age aspect.
 Recent research Duan et al.
 (2024) publishes an evaluation for well-known LLMs on age bias through 50 multi-choice questions; unlike it focuses on discriminatory narratives towards specific age groups, our investigation is running at an implicit level.
-We argue that understanding the underlying value systems is crucial, as the value discrepancies between users and LLMs can significantly impact their adoption of LLMs, even though the explicit discrimination is rectified, as exemplified in our discussion on technology attitudes in Sec 1.
-Figure 2: Alignment rank of values of LLMs over different age groups in specific Countries.
-See results on more models and countries in Appx E and F .
-Rank 1 on an specific age group represents that this age group has the narrowest gap with LLM in values.
-A increasing monoticity indicates a closer alignment towards younger groups.
+We argue that understanding the underlying value systems is crucial, as the value discrepancies between users and LLMs can significantly impact their adoption of LLMs, even when the explicit discrimination is rectified, as exemplified in technology attitudes discussed in Sec 1.
 3 Analytic Method
@@
 Dataset.
-We derive human values utilizing the 7th wave of the World Values Survey (WVS) Haerpfer et al.
+We derive human values utilizing a well-established survey dataset, the 7th wave of the World Values Survey (WVS) Haerpfer et al.
 (2020).
 The survey systematically probes 94k individuals globally on 13 categories, covering a range of social, political, economic, religious, and cultural values.
-See an introduction of WVS in Appx A.
+See more about WVS in Appx A.
 Each inquiry is a single-choice question.
-Responses are numeric, quantifying the inclination on the options, e.g., “1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree”.
-Negative number is possible for coding exceptions such as “I don’t know”.
-To assess human values, we group the respondents by age group  [^1] Age groups are recorded as 18-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country.
+Responses are numeric, quantifying the inclination on the options, e.g., “1:Strongly agree, 2:Agree, 3:Disagree, 4:Strongly disagree".
+Negative number is possible for coding exceptions such as “I don’t know".
+To assess human values, we group the respondents by age group 11118-24, 25-34, 35-44, 45-54, 55-64, and 65+ and country.
 Subsequently, we compute the average values for each age group and country to represent their respective cohorts, ignoring the invalid negative numbers.
@@
 ChatGPT(GPT-3.5-turbo 0613)
-InstructGPT (GPT-3.5-turbo-instruct)
-Mistral (mistral-7B-v0.1)
-Vicuna (vicuna-7b-v1.5)
-FLAN-T5 (flan-t5-xxl)
-FLAN-UL2 (flan-ul2)
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
+[…]
+*[Uncaptioned image] 0
 Table 1: Model description.
+[…]
+*[Uncaptioned image] 0
 : commercial models,
+[…]
+*[Uncaptioned image] 0
 : open models,
+[…]
+*[Uncaptioned image] 0
 : chat-based,
+[…]
+*[Uncaptioned image] 0
 : completion-based,
+[…]
+*[Uncaptioned image] 0
 : RLHF, and
+[…]
+*[Uncaptioned image] 0
 : training with instructions.
+Figure 2: Alignment rank of values of LLMs over different age groups in specific Countries.
+See results on more models and countries in Appendix D and E .
+Rank 1 on a specific age group means that this age group has the narrowest gap with LLM in values.
+An increasing monoticity indicates a closer alignment towards younger groups.
 Prompts.
@@
 Eventually, we build a set of eight distinct prompts per inquiry.
-Please see prompt design details in Tab 3 in Appendix.
-Through a careful analysis on the prompt responses (Appx B), We observe unstableness of LLM’s responses to prompt variations.
-However, multiple prompt trials assists with achieving a convergence point.
+Please see prompt design details in Tab 8.
+Through a careful analysis of the prompt responses (Appx B), we observe the unstableness of LLM’s responses to prompt variations.
+However, multiple prompt trials assist with achieving a convergence point.
 On 95.5% of questions, more than half of the eight prompts led to responses centered on the same choice or adjacent options, and thus we believe it is acceptable to consider the average of the outcomes across the eight prompt variations as the LLM’s final responses to WVS.
-In addition, due to the instability of LLMs in following instructions, we encountered seven types of unexpected reply and present our coping methods for each, as summarized in Tab 4.
+In addition, due to the instability of LLMs in following instructions, we summarize seven types of unexpected replies and present our coping methods for each in Tab 3.
 In the process of averaging responses, we ignore the invalid negative numbers, as we did in calculating human values.
-For reproducing our work, parameter setting and prompting details are reported in Appendix D.
+For reproducing our work, prompting details are reported in Appx C.
 3.3 Measures
@@
 Note the acquisition of numeric responses for human groups and LLM has been illustrated in Sec 3.1 and 3.2.
-By collecting 372 value vectors that represent people across 62 countries and 6 age groups, along with a value vector for the LLM to compare, we utilize principle component analysis (PCA) Tipping and Bishop (1999) on totally 373 value vectors for representation learning.
+By collecting 372 value vectors that represent people across 62 countries and 6 age groups, along with a value vector for the LLM to compare, we perform min-max normalization, normal standardization, and then conduct principle component analysis (PCA) Tipping and Bishop (1999) on a total of 373 value vectors for representation learning.
 We acquire value representations for all groups with the dimensionality of three.
-Our consideration of using PCA is added in Appx C.
-[…]
-Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value representation for the $i$th age group be $x_{c,i},y_{c,i},z_{c,i}]$.
+Our consideration of using PCA is in Appx G.1.
+[$x_{c},y_{c},z_{c}]=PCA\_transform([r_{1},r_{2},...r_{n_{c}}])$
+Let $i$ be the index of age group in [18-24, 25-34, 35-44, 45-54, 55-64, 65+] and the value representation for the $i$th age group be $[x_{c,i},y_{c,i},z_{c,i}]$.
 We derive three metrics below for our further analyses:
@@
 […]
-Trend Coefficient, the slope of the value gap between LLM and humans across six age groups. $\alpha$ is the slope we would like to fit by linear regression.
+Trend Coefficient, the slope of the value gap between LLM and humans across six age groups. Let $\alpha_{c}^{*}$ be the optimal coefficient to fit the linear relation:
 […]
 […]
+Our reasons for these measure designs are detailed in the Appx G.
 4 Aligning with Which Age on Which Values?
 Trend Observation.
-As shown in Fig 1, we observe a general inclination of popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient.
-Fig 2 exemplifies the bias for LMMs across six age groups in several countries.
-Due to the limited paper pages, results on other LLMs and countries can be found in Appx E and F.
-Significant testing procedure is available in Appx G.
-We observe that in the US and China, as countries of large population, the models tend to have a higher alignment rank on younger groups on the most categories, despite few exceptions (e.g., happiness and well-being).
+Fig 2 exemplifies the bias for LLMs across six age groups in several countries.
+Due to the limited paper pages, results on other LLMs and countries can be found in Appx D and E.
+As it is not intuitive to see a bias towards younger people in these decoupled results, we summarize the performance of all LLMs in the US, as shown in Fig 1.
+Then we observe a general inclination of popular LLMs favoring the values of younger demographics in the US on different value categories, indicated by the trend coefficient.
+Significance testing procedure is available in Appx F.
+We observe that in the US and China, as countries with large populations, the models tend to have a higher alignment rank on younger groups on most categories, despite few exceptions (e.g., happiness and well-being).
 However, in Ethiopia and Nigeria (Tab 15), the inclination is less evident.
 We leave this phenomenon for future study.
+Case Study.
+In Fig 3, we show two representative prompts and their responses from ChatGPT and human groups, to exemplify values where ChatGPT displays a clear inclination toward a specific age group.
+Note LLM values can be far away from all human age groups, as depicted in the second sub-figure.
+We discuss this point in Appx G.2.
 Figure 3: Two WVS prompts and their responses from LLMs and humans (in purple).
-Case Study.
-In Fig 3, we show two representative prompts and their responses from ChatGPT and human groups, to illustrate sample values where ChatGPT exhibits a clear bias toward a specific age group.
 5 The Effect of Adding Identity in Prompts
 Prompt Adjustment.
-To analyze if adding age identity in the prompt helps to align values of LLM with the targeted age groups, we adjust our prompts by adding a sentence like “Suppose you are from [country] and your age is between [lowerbound] and [upperbound].” at the beginning of the required component of the original prompt and get responses that corresponds with six age groups.
+To analyze if adding age identity in the prompt helps to align values of LLM with the targeted age groups, we adjust our prompts by adding a sentence like “Suppose you are from [country] and your age is between [lowerbound] and [upperbound].” at the beginning of the required component of the original prompt and get responses that correspond with six age groups.
 Observation on Gap Change.
 We illustrate the change of Euclidean distance between values of LLM and different age groups after adding identity information.
-As is presented in Fig 4, in eight out of thirteen categories (No.1,2,4,5,7,8,9,12) no improvement is observed.
+As is presented in Fig 4, in eight out of thirteen categories (No.1,2,4,5,7,8,11,12) no improvement is observed.
 Figure 4: Change of Euclidean distance after adding identity information. The compared data is from values of ChatGPT and humans from different age groups in the US.
@@
 We also showcase a successful calibration example for a question about the source of acquiring information in Fig 5.
-The value pyramid illustrates LLMs’ responses for different age ranges compares to the answers from the U.S.
+The value pyramid illustrates LLMs’ responses for different age ranges compared to the answers from the U.S.
 population.
@@
 Figure 5: Value Pyramid of U.S population (left) and ChatGPT (right) for an inquiry on the frequency of using radio news.
-6 Recommendations for Future Work
-We have observed that simply including an age in prompts fails to eliminate the value disparity for the targeted age groups.
-Out of the thirteen categories inquired upon, eight have shown no improvement.
-To this end, we recommend a careful data curation during pretraining.
-Doing so involves a deliberate and thoughtful selection of data sources that are diverse and representative of various age groups.
-By doing so, we can ensure that the model’s training material reflects a wide range of perspectives and experiences, thereby reducing biases and disparities in the model’s responses.
-We also recommend a consideration of human feedback optimization (e.g., RLHF).
-Through this iterative process, LLMs can learn to generate responses that fit better with the needs of different age groups.
-These strategies help mitigate the value disparities associated with targeted age groups, enhancing the LLM’s abilities to be more equitable and inclusive.
+6 Further Discussion on the Age Bias Observed in LLMs
+In this study, we have shown how LLMs are not representative of the value systems of older adults.
+Although further validation is necessary for a solid conclusion, we believe there may be several potential harms arising from this bias:
+  • Older adults tend to place greater trust in established organizations, particularly when it comes to security concerns (as illustrated in Fig 1).
+An LLM unaware of these differences may pose greater risks to older users, who may be less prepared to identify misinformation from what appears to be a credible source (e.g., LLM itself).
+This could amplify the harm caused by LLM-generated hallucinations when letting LLMs serve aged people.
+  • LLMs may offer less empathetic interactions to older adults by failing to account for their traditional beliefs, leading to less respectful exchanges.
+  • For older adults, who are often less inclined towards new technologies, interacting with LLMs embodying the values of tech-savvy users could further alienate them.
+As shown in Fig 3, many older adults still rely on the radio for news, while younger people predominantly use the internet.
+7 Suggestions on Age-aware Alignment for Future Work
+Although we have shown that LLMs are not representative of the value systems of older adults, our study is not intended to promote a naive copy of the values of different age groups to achieve alignment.
+Simplistically applying statistical knowledge of the values of a particular age group might reinforce stereotypes rather than promote genuine alignment.
+For example, consider whether LLMs should adopt the value that the older generation is less tech-savvy and thus develop the stereotype that an older user would primarily obtain news from the radio rather than social media.
+However, as illustrated in Fig 6, while fewer older adults rely on social media for information, a significant portion still does.
+Therefore, LLMs must be aware of statistical discrepancies but should avoid brute-force applying statistics to any individual, as a brute-force application often only considers the mean instead of other qualities, such as variance, outliers, and so on.
+Thus, to facilitate a true age-aware alignment, we recommend researchers to rely on the following rules of thumb:
+  • Avoid naively applying statistical knowledge of the values of a particular age group, as this can reinforce stereotypes instead of promoting genuine alignment.
+  • Develop strategies that promote true age-sensitive interactions, emphasizing age-aware helpfulness and harmlessness, grounded in an understanding of value discrepancies across generations.
+Figure 6: People’s preference on obtaining information from social media across different age groups in the US population
+Achieving age-aware alignment requires LLMs to be sensitive to value differences across age groups and to build on these insights to offer helpful and harmless responses.
+For example, when engaging with older users, instead of brute-force assuming they are lagging behind new technology, a well-aligned system should keep tracking their understanding of the ongoing topics, offering more detailed explanations and minimizing the use of neologisms only when confusion arises.
+To achieve such age-sensitive interactions, exploring an effective feedback-acquiring method during interactions that complies with the real age-tailored connotation of helpfulness and harmlessness is meaningful.
+Although challenging, we believe this is a vital direction for future research.
 8 Conclusion
@@
 Our findings suggest a general inclination of LLM values towards younger demographics.
-Our study contributes to raising attention to the potential age bias in LLMs and advocate continued efforts from the community to address this issue.
-Moving forward, efforts to calibrate value inclinations in LLMs should consider the complexities involved in prompting engineering and strive for equitable representation across diverse age cohorts.
+Our study contributes to raising attention to the potential age bias in LLMs and advocates continued efforts from the community to address this issue.
+Moving forward, efforts to calibrate value inclinations in LLMs should consider the complexities involved in prompt engineering and strive for equitable representation across diverse age groups.
 Limitations
 There are several limitations in our paper.
-Firstly, due to the time and cost, we were not able to try more sophisticated prompts for the age alignment, which may effectively eliminate the value disparity with targeted age groups.
-Secondly, our analysis relies on the questionnaire of WVS.
-However, their question design is not perfectly tailored for characterizing age discrepancies, which limits the depth of sights we could get from analysis.
-Finally, the range of LLMs in our analysis could be expanded.
+Firstly, Fig 3 may raise questions concerning the importance of any trends in light of LLM values not resembling any age group of humans.
+We conjecture that due to the nature of Human Preference Optimization Rafailov et al.
+(2024); Ouyang et al.
+(2022), LLMs develop extreme preferences (e.g., manifest an extreme atheist).
+The resulting LLMs will thus be unlike the subtler preferences of humans.
+Our study does not focus on the absolute difference between LLMs and humans, but instead emphasizes the inclination, as we have explained in Appendix G.2.
+However, future work is needed to reflect on the current process of Human Preference Optimization, especially on whether it will be problematic or acceptable if we over-align LLMs with human preference.
+Secondly, due to time and cost considerations, we were not able to try more sophisticated prompts for age alignment, which may effectively eliminate the value disparity with targeted age groups.
+Finally, our analysis relies on the questionnaire of WVS.
+However, their question design is not perfectly tailored for characterizing age discrepancies, which limits the depth of sight we could get from analysis.
 Ethics Statement
-Several ethical considerations have been included thorough our projects.
-Firstly, the acquisition of WVS data is under the permission of data publisher.
-Secondly, we carefully present our data analysis results with an academic honesty.
-This project is under a collaboration, we well-acknowledge the work of each contributor and ensure a transparent and ethical process thorough the whole collaboration.
-Finally, we leverage the ability of AI-assistants to help with improving paper writing while we guarantee the originality of paper content and have reviewed the paper by every word.
+Several ethical considerations have been included through our projects.
+Firstly, the acquisition of WVS data is under the permission of the data publisher.
+Secondly, we carefully present our data analysis results with academic honesty.
+This project is under a collaboration, we well-acknowledge the work of each contributor and ensure a transparent and ethical process throughout the whole collaboration.
+Finally, we leverage the ability of AI assistants to help with improving paper writing while we guarantee the originality of paper content and have reviewed the paper by every word.
+Acknowledgements
+We thank the anonymous reviewers for their constructive feedback, and the members of the Language and Information Technologies lab at the University of Michigan for the insightful discussions during the early stage of the project.
+We thank Trenton Chang for his insightful questions in AI snippet activity at the University of Michigan.
+This project was partially funded by a National Science Foundation award (#2306372) and a grant from OpenAI.
+Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation or OpenAI.
 Appendix A World Value Survey
-The WVS [^3] https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values.
+The WVS333 The data can be downloaded via https://www.worldvaluessurvey.org/wvs.jsp survey is conducted every five years, which systematically probes individuals globally on social, political, economic, religious, and cultural values.
 We share a page of WVS questionnaire in Tab 7.
 See the statistics of inquiries in Fig 2.
-Demographic statistics of WVS is accessible via Document-Online analysis.
-Note that we remove ten of them that requires demographic information, as these are impossible for applying to an LLM lacking demographic data, and keep 249 inquiries as our final choices for prompting.
+Demographic statistics of WVS are accessible via Document-Online analysis.
+Note that we removed ten of them that require demographic information, as these are impossible to apply to an LLM lacking demographic data, and kept 249 inquiries as our final choices for prompting.
 Value Category # Inquiry Example
@@
 (c) Example Prompt
-Table 3: Prompt Pipeline Details
+Figure 8: Prompt Pipeline Details
 Unexpected Reply Type Example Coping Method
-returning null value { ”Q1”: null} map null into missing code -2
+returning null value { "Q1": null} map null into missing code -2
 unprompted responses answer Q1 to Qn when only asking Qn-m to Qn keep the answers of asked questions
-redundant texts ”Answer = {‘Q1’, 1}” extract the json result
+redundant texts "Answer = {‘Q1’, 1}" extract the json result
 substandard json Q1:‘1’ manually correct
@@
     Despite potential variations in answers induced by prompt variation, we found for 95.5% of inquiries, more than half of the responses are centered on the same choice or its adjacent options.
-Adjacent option is a score equal to the majority score +/- 1.
+The adjacent option is a score equal to the majority score +/- 1.
 Eventually, while discovering the unstableness of LLM outputs, we believe it is reasonable to use the average score from eight prompts as a representative value.
-Appendix C Reasons of Applying PCA
-  1. 1.
-    Each question in WVS ought not to be equally important.
-Furthermore, for the questions belonging to a certain category, they correlate with each other.
-To this end, we need to find out the principle components among multi inquiries.
-  2. 2.
-    PCA here is also used as an unsupervised representation learning method.
-Compared to utilizing original data, the representations learned from hundreds of comparable examples (372 value vectors from different country and age groups) will mitigate the curse of dimensionality and other undesired properties of high-dimensional spaces.
-Other representation learning methods are also applicable.
-As the medium number of original dimensionality for all categories is 11, PCA is enough to handle the learning problem.
-Appendix D Prompting Details
+Appendix C Prompting Details
 Our prompting process can be described as three steps below:
@@
 For the deployment of open-sourced models, we ran either model on a single A40 GPU with float16 precision.
-When prompting, we prompt models with a temperature 1.0, max token length 1024, random seed 42.
-Appendix E Results on Other LLMs
-In the section, we supplement the alignment ranking results on InstructGPT (Fig 7), FLAN-T5-XXL (Fig 8) and FLAN-UL2 (Fig 9) respectively.
-Appendix F Results on Other Countries
-We have extended our analysis to include alignment results from an additional four pairs of countries: Argentina and Brazil (Fig 7), Ethiopia and Nigeria (Fig 8), Germany and Great Britain (Fig 9), and Indonesia and Malaysia (Fig 10).
-Appendix G Significant Test
-In this section, we conduct two kinds of significant tests to support our study: (1) we use MANOVA to test the significant difference among human values from different age groups, and (2) we use t-distribution to test the significant tendency of LLMs towards younger group.
-Notes our focus lies in characterizing the inclination of LLM values towards specific age groups.
-So, we are claiming significant tendency instead of claiming LLMs significantly assemble of any specific age groups.
-G.1 Significant Test for the Discrepancy among Human Age Groups
-Our analysis should base on a reasonable precondition that in WVS, human values significantly diverse cross different age groups.
-We used MANOVA (multivariate analysis of variance) to test the significant difference of human values cross all age groups, as shown below:
+When prompting, we prompt models with a temperature 1.0, max token length 1024, and random seed 42.
+Appendix D Results on Other LLMs
+In the section, we supplement the alignment ranking results on InstructGPT (Fig 9), FLAN-T5-XXL (Fig 10) and FLAN-UL2 (Fig 11), Mistral (Fig 12) and Vicuna (Fig 13) respectively.
+Appendix E Results on Other Countries
+We have extended our analysis to include alignment results from an additional four pairs of countries: Argentina and Brazil (Tab 14), Ethiopia and Nigeria (Tab 15), Germany and Great Britain (Tab 16), and Indonesia and Malaysia (Tab 17).
+Appendix F Significance Test
+In this section, we conduct two kinds of significance tests to support our study: (1) we use MANOVA to test the significant difference among human values from different age groups, and (2) we use t-distribution to test the significant tendency of LLMs towards younger groups.
+Notes our focus lies in characterizing the inclination of LLM values toward specific age groups.
+That is to say, we are claiming a significant tendency over age, rather than claiming LLMs significantly resemble any specific age group.
+We make a deeper discussion about our declaration in the section on Limitations.
+F.1 Significance Test for the Discrepancy among Human Age Groups
+Our analysis should be based on a reasonable precondition that in WVS, human values are significantly diverse across different age groups.
+We used MANOVA (multivariate analysis of variance) to test the significant difference in human values across all age groups, as shown below:
 Null hypothesis ($H_{0}$): the age group has no effect on any responses to the survey questions
 Statistics: Wilks’ lambda
-Result: See Tab 5. In conclusion: We reject the null hypothesis with p-value ¡ 1e-4
-Country Value Num DF Den DF F Value Pr ¿ F (p-value)
+Result: See Tab 4. In conclusion: We reject the null hypothesis with p-value < 1e-4
+Country Value Num DF Den DF F Value Pr > F (p-value)
 US 0.07 176.00 1631.00 124.82 0.0000*
@@
 Nigeria 0.13 176.00 614.00 23.18 0.0000*
-Table 5: P-values of value difference among different age groups in specific countries. * indicates p-value¡1e-4
-G.2 Significant Test for Trend Coefficient
+Table 4: P-values of value difference among different age groups in specific countries. * indicates p-value<1e-4
+F.2 Significance Test for Trend Coefficient
 As it may be hard to interpret the trend coefficient in Fig 1 on some categories (e.g., perception of corruption).
@@
 Results: see Tab 5.
+Appendix G Our Consideration on Measure Design
+G.1 Reasons for Applying PCA
+We choose PCA for the following reasons:
+  1. 1.
+    Each question in WVS ought not to be equally important.
+Furthermore, for the questions belonging to a certain category, they correlate with each other.
+To this end, we need to find out the principal components among multiple inquiries.
+  2. 2.
+    PCA here is also used as an unsupervised representation learning method.
+Compared to utilizing original data, the representations learned from hundreds of comparable examples (372 value vectors from different countries and age groups) will mitigate the curse of dimensionality and other undesired properties of high-dimensional spaces.
+Other representation learning methods are also applicable.
+As the medium number of original dimensionality for all categories is 11, PCA is enough to handle the learning problem.
+Furthermore, we set the target number of PCA components to three.
+We empirically set so, considering the medium number of original dimensionality for all categories is eleven.
+Then we validate this parameter by calculating the percentage of variance explained by each of the selected components.
+If all components are stored, the sum of the ratios is equal to 1.0.
+The explained variance ratio of keeping three dimensions is an average of no less than 0.72 in all categories of six models, which we believe is acceptable.
+G.2 Consideration of Using the Rank of Difference as Measurement
+In Sec 3.3, we utilize the rank of difference to characterize the value discrepancies and the trend coefficient over age.
+Presenting rank is simple and convenient for data visualization.
+However, using the rank of difference may ignore the magnitude (the absolute value) of difference that is (1) among the different age groups of humans or (2) between LLM values and specific age groups of humans.
+We further clarify that:
+(1) Appx F.1 has shown significant value discrepancies among different age groups of humans in the countries we experiment on.
+So, using the rank of difference would not exaggerate a significant disparity between human age groups to observe, as the discrepancies have existed significantly.
+(2) As shown in the second sub-figure of Fig 3, it is possible that LLMs values are far away from all human age groups.
+Such discrepancies also would not reflect on the rank of difference.
+However, our study focus lies in characterizing the inclination of LLM values towards specific age groups.
+That is to say, we are claiming a significant tendency over age, rather than claiming LLMs significantly resemble any specific age group.
+We make a deeper discussion about our declaration in the section of Limitations.
 Category ChatGPT InstructGPT Mistral Vicuna Flan-t5 Flan-ul
@@
 Political Culture and Political Regimes 0.000* 0.000* 0.000* 0.005* 0.957 0.872
-Table 6: P-values of trend coefficients for each model on each value category. * indicates p-value¡0.1
+Table 5: P-values of trend coefficients for each model on each value category. * indicates p-value<0.1
 Figure 9: Alignment rank of values of InstructGPT over different age groups in the US.
-Rank 1 on an specific age group represents that this age group has the narrowest gap with InstructGPT in values.
-A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
+Rank 1 on a specific age group represents that this age group has the narrowest gap with InstructGPT in values.
+An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
 Figure 10: Alignment rank of values of FLAN-T5-XXL over different age groups in the US.
-Rank 1 on an specific age group represents that this age group has the narrowest gap with FLAN-T5-XXL in values.
-A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
+Rank 1 on a specific age group represents that this age group has the narrowest gap with FLAN-T5-XXL in values.
+An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
 Figure 11: Alignment rank of values of FLAN-UL2 over different age groups in the US.
-Rank 1 on an specific age group represents that this age group has the narrowest gap with FLAN-UL2 in values.
-A increasing monoticity indicates a closer alignment towards younger groups, vice versa.
-Table 7: Alignment rank of LLMs over different age groups in Argentina and Brazil. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
-Table 8: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria. LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
-Table 9: Alignment rank of LLMs over different age groups in Gemany and Great Britain.
+Rank 1 on a specific age group represents that this age group has the narrowest gap with FLAN-UL2 in values.
+An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
+Figure 12: Alignment rank of values of Mistral over different age groups in the US.
+Rank 1 on a specific age group represents that this age group has the narrowest gap with Mistral in values.
+An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
+Figure 13: Alignment rank of values of Vicuna over different age groups in the US.
+Rank 1 on a specific age group represents that this age group has the narrowest gap with Vicuna in values.
+An increasing monoticity indicates a closer alignment towards younger groups, vice versa.
+Figure 14: Alignment rank of LLMs over different age groups in Argentina and Brazil.
 LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
-Table 10: Alignment rank of LLMs over different age groups in Indonesia and Malaysia.
+Figure 15: Alignment rank of LLMs over different age groups in Ethiopia and Nigeria.
 LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
+Figure 16: Alignment rank of LLMs over different age groups in Gemany and Great Britain.
+LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.
+Figure 17: Alignment rank of LLMs over different age groups in Indonesia and Malaysia.
+LLM tested in each image is (a) ChatGPT, (b) InstructGPT, (c) Mistral, (d) Vicuna, (e) Flan-t5-xxl, and (f) Flan-ul.